OpenStudy (anonymous):

Solve for x: 2 over quantity x plus 2 plus 1 over 5 equals 6 over quantity x plus 5

OpenStudy (anonymous):

\[\frac{ 2 }{ x+2 }+ \frac{ 1 }{ 5 }= \frac{ 6 }{ x+5 }\]